OpenStudy (anonymous):

find k such that the line is tangent to the graph of the function fuction= f(x)=x^2-kx line y=4x-9

OpenStudy (campbell_st):

find the 1st derivative..

OpenStudy (anonymous):

i did

OpenStudy (campbell_st):

ok... the gradient/slope of the given line y = 4x - 9 is..?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ok what do i do

OpenStudy (campbell_st):

so you need to set the 1st derivative equal to 4 and solve for k

OpenStudy (anonymous):

i solved for k k=2x-4

OpenStudy (campbell_st):

ok... so you have \[x^2 - (2x - 4)x\] you need to now find the point of intersection so equate the parabola and line x^2 - (2x - 4)x = 4x - 9 after simplifying -x^2 = -9 x^2 = 9 so x = 3 (3, 3) is the point of intersection.

OpenStudy (anonymous):

thanks!

OpenStudy (campbell_st):

so m = 4 when x = 3 so using the 1st derivative 4 = 2(3) - k k = 2 hope it makes sense... it was a bit more difficult than I thought...